if one car leaves as 11:45 and travels at 30 miles per hour and anothere car leaves at 12:00 and travels at 45 miles per hour what time will they meet
Answer by rfer(16322)

30/4=7.5 mi lead
45-30=15 mph overtake speed
7.5/15=0.5 hrs
12:00+00:30=12:30 meet time